Numerical Analysis I, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 10

1. Section 4.7: Problems 1(f), 2(f), 5, 7, 8.

2. Section 4.9: Problem 4.

Remark: The change of variable t = 1/(x+1) probably works as well (without splitting the domain into two parts). You are encouraged to explore with this approach.